OpenStudy (anonymous):

Lets see here... \[f(x)=\frac{x^2}{(1-2x)^2}\] \[=x^2\frac{d}{dx}\frac{1}{(1-2x)}\] \[=x^2\frac{d}{dx}\left(\sum_{n=0}^{\infty}(2x)^n\right)\] \[=x^2\sum_{n=0}^{\infty}2nx^{n-1}\] \[=\sum_{n=0}^{\infty}2nx^{n+1}\]

OpenStudy (anonymous):

how did I do

OpenStudy (anonymous):

a little mistake in derivating...

OpenStudy (anonymous):

pretty sure u missed a coeff

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

multiply by -.5 and u should be good!

OpenStudy (anonymous):

and we must have the condition \(|2x|<1\)

OpenStudy (anonymous):

good catch.

OpenStudy (anonymous):

actually not just a coefficient...

OpenStudy (anonymous):

lol yeah her second differentiation is wrong too...

OpenStudy (anonymous):

and when you fix that, I don't think it will simplify nicely.

OpenStudy (anonymous):

@MathSofiya note that \[\frac{d}{dx} (2x)^n=2n(2x)^{n-1} \]

OpenStudy (anonymous):

or at least not without havving a clunky 2^n

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}2n(2x)^{n-1}\] what should I do about the x^2?

OpenStudy (anonymous):

well as i said if u distribute the n-1 exp over the 2 and the x

OpenStudy (anonymous):

then u can have a 2^n term and ur x^n+1 term

OpenStudy (anonymous):

\[f(x)=\frac{x^2}{(1-2x)^2}\] \[=\frac{x^2}{2}\frac{d}{dx}\frac{1}{(1-2x)}\] \[=\frac{x^2}{2}\frac{d}{dx}\left(\sum_{n=0}^{\infty}(2x)^n\right) \ \ \ \ if \ \ \ |2x|<1\] \[=\frac{x^2}{2}\sum_{n=0}^{\infty}2n(2x)^{n-1}\] \[=\sum_{n=0}^{\infty}n2^{n-1}x^{n+1}\]

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}2n2^{n-1}x^{n+1}\]

OpenStudy (anonymous):

well yes but as i said u are missing a -.5 and the 2 times 2^n-1 is just 2^n

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yeah np

OpenStudy (anonymous):